You are looking at producing a 3D model of a solid shape you want to manufacture. The shape is a cone which is bounded by the line 𝑦=1/2π‘₯ rotated about the π‘₯βˆ’π‘Žπ‘₯𝑖𝑠 between π‘₯=0 π‘Žπ‘›π‘‘ π‘₯=5?


Use calculus to find its volume.


1
Expert's answer
2021-06-01T17:40:59-0400

Solution

"V=\pi\int_{0}^{5}{y^2dx}=\pi\int_{0}^{5}[x/2]^2dx =\frac{π}{4} \frac{x^3}{3}|_0^5=125π/12 = 32.725"